OpenStudy (anonymous):

solve (343)^1/3 A. 49 B. 7 C. 21 D. 19

OpenStudy (cwrw238):

just use your calculator to do this either 343 to the power (1/3) or the cube root

OpenStudy (cwrw238):

or without a calculator cube each of the options and see which one fits

OpenStudy (anonymous):

so I would have to multiply a,b,c or d till I get 343?

OpenStudy (anonymous):

thats is one way you could try

OpenStudy (cwrw238):

try each in turn - you know what cube it means? / multiplying a number by itself 3 times eg 4 * 4 * 4

OpenStudy (cwrw238):

10*10*10 = 1000 - now you should see which option is correct

OpenStudy (anonymous):

Is it B?

OpenStudy (cwrw238):

yea

OpenStudy (anonymous):

Yay! Thank you :)

OpenStudy (cwrw238):

7*7*7 = 49*7 = 343

OpenStudy (anonymous):

ok 1 more thing about the "1/3" if it was like "1/4" or "1/5" would I multiply each number 4 or 5 times or just till I get the number that matches up with the "343" place?